Two trains leave stations
494 miles apart at the same time and travel toward each other. One train travels at
105 miles per hour while the other travels at
85 miles per hour. How long will it take for the two trains to meet?
Do not do any rounding.

Answer by stanbon(75887)

Two trains leave stations
494 miles apart at the same time and travel toward each other. One train travels at
105 miles per hour while the other travels at
85 miles per hour. How long will it take for the two trains to meet?
Do not do any rounding.
------
Faster train DATA:
distance = x miles ; rate = 105 mph ; time = x/105 hrs
----
Slower train DATA:
distance = 494-x miles ; rate = 85 mph ; time = (494-x)/85 hrs
----------------------
Equation:
time = time
x/105 = (494-x)/85
85x = 105*494-105x
-----
190x = 105*494
x = 273 miles
---
Ans:: time = 273/105 = 2.6 hrs = 2hrs 36min